40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3937     Entry time: Wed Nov 17 02:53:41 2010
Author: yuta 
Type: Update 
Category: IOO 
Subject: placed new PRM to BS table 

(Kiwamu, Yuta)

Background:
  Yesterday, we aligned the Faraday and the beam reached SM2 at BS table.
  Today, we placed a new PRM tower to BS table.

What we did:

  1. Moved IPPO, IPPOSSM1, IPPOSSM3, IPANGSM1, IPANGSM2 out from the BS chamber.

  2. Moved SRM tower(at PRM's place) to the ITMX chamber.

  3. Placed the new PRM tower at the BS chamber.

  4. Adjusted positions of the OSEMs for PRM and BS so that the sensor output can have roughly half of their maximum.

  5. Checked damping servo for PRM and BS. They were working and helped us when adjusting OSEM positions.

  6. Placed IPPO back and using SM2, made the beam hit PR2 at ITMX table.

  7. Aligned the PRM so that the reflected beam path overlaps the incident beam.
     We checked it by looking at MMT1.
     For the alignment, we used IFO align sliders(C1:SUS-PRM_PIT_COMM, YAW_COMM).
        To use them, we rebooted c1susaux.

Result:
  1. The new PRM tower is placed.

  2. OSEM sensor outputs for PRM and BS are;

(V) PRM BS
max current value max current value
ULSEN 1.72 1.006 1.50 0.757
URSEN 1.66 0.918 1.57 0.821
LRSEN 1.92 1.304 1.57 0.821
LLSEN 2.06 1.031 1.38 0.704
SDSEN 9.21 4.366 1.57 0.821

    We changed PRM aligning slider values, and they changed OSEM sensor outputs. We set the slider values to 0 when adjusting OSEM positions.

ELOG V3.1.3-